OpenStudy (anonymous):

Find the slope of a line perpendicular to a line through the given points. E(5, 7), F(3 , 1)

OpenStudy (amistre64):

E-F is the key

OpenStudy (amistre64):

or F-E ... but it turns out the same

OpenStudy (amistre64):

y/x = parallel -x/y = perp

OpenStudy (amistre64):

E(5, 7) -F(3 , 1) -------- x , y can you tell me the values i get for x and y?

OpenStudy (anonymous):

1= Slope of the line passing through the given points= 7-1/5-3 =6/2 = 3..........................1 m2 = slope of line perpendicular to the given line....................................… m1*m2 = -1 3*m2 = -1 m2 = -1/3....................................…

OpenStudy (anonymous):

So it's -1/3?

OpenStudy (amistre64):

if you would participate more, it might help you out ....

OpenStudy (anonymous):

Pardon?

OpenStudy (anonymous):

yerp

OpenStudy (anonymous):

Thanks Law!

OpenStudy (anonymous):

np